la.) TTM: 15 years. YTM: 7%. C: 7%. PV: 2 1b.) TTM: 10 years. YTM: 7%. C: 7%. PV:? Ic.) TTM:5 years. YTM: 7%. C: 7%. PV:? 2a.) TTM: 15 years. YTM: 9% C: 7%. PV:? 2b.) TTM: 10 years. YTM: 9%. C: 7%. PV:7 2c.) TTM: 5 years. YTM: 9%. C: 7%. PV: ? 3a.) TTM: 15 years. YTM: 5%. C: 7%. PV:7 3b.) TTM: 10 years. YTM: 5%. C: 7%. PV: ? 3c.) TTM: 5 years. YTM: 5%. C: 7%. PV: 7
Questions for Problems 1 3 a.) What happens to the price of par, discount, and premium bonds as the remaining maturity declines? Note that the YTM remains unchanged. Only the remaining maturity changes. b,) Explain why prices move this way. IT
Answer #1

1a)

A Par value (FV) 1,000 7.00% 1 70 15 4 DAxB/C Interest per period (PMT) E Number of years to maturity Number of compounding p

1b)

A Par value (FV BlCoupon rate ClNumber of compounding periods per vear 1,000 7.00% 4 | D = AxB/C| Interest per period (PMT 70 10 El Number of vears to maturit Number of compounding periods till maturity 10 7.00% GMarket rate of return/Required rate of return Market rate of return/Required rate of return per G/C period (RATE) 700% PV(RATE,NPER,PMT, FV)*-1 Price of the bond (PV 10 Price of the bond (PV) 1,000.00 12 =PV(796, 10,70, 1000)-1

1c)

A Par value (FV BlCoupon rate ClNumber of compounding periods per vear 1,000 7.00% 4 | D = AxB/C| Interest per period (PMT 70 El Number of vears to maturit Number of compounding periods till maturity GMarket rate of return/Required rate of return 7.00% Market rate of return/Required rate of return per G/C period (RATE) 700% PV(RATE,NPER,PMT, FV)*-1 Price of the bond (PV 10 Price of the bond (PV) 1,000.00 12 =PV( 796,5,70, 1000)-1

2a)

A Par value (FV BlCoupon rate ClNumber of compounding periods per vear 1,000 7.00% 4 | D AXB/C| Interest per period (PMT 70 E